# AnomalyControl: Learning Cross-modal Semantic Features for Controllable Anomaly Synthesis

### v1

@@ Abstract
Anomaly synthesis is a crucial approach to augment abnormal data for advancing anomaly inspection. Based on the knowledge from the large-scale pre-training, existing text-to-image anomaly synthesis methods predominantly focus on textual information or coarse-aligned visual features to guide the entire generation process. However, these methods often lack sufficient descriptors to capture the complicated characteristics of realistic anomalies (e.g., the fine-grained visual pattern of anomalies), limiting the realism and generalization of the generation process. To this end, we propose a novel anomaly synthesis framework called AnomalyControl to learn cross-modal semantic features as guidance signals, which could encode the generalized anomaly cues from text-image reference prompts and improve the realism of synthesized abnormal samples. Specifically, AnomalyControl adopts a flexible and non-matching prompt pair (i.e., a text-image reference prompt and a targeted text prompt)  [^1] Non-matching: the text-image reference prompt and targeted text prompt are only required to describe the same anomaly type without needing to match in surface details or materials., where a Cross-modal Semantic Modeling (CSM) module is designed to extract cross-modal semantic features from the textual and visual descriptors. Then, an Anomaly-Semantic Enhanced Attention (ASEA) mechanism is formulated to allow CSM to focus on the specific visual patterns of the anomaly, thus enhancing the realism and contextual relevance of the generated anomaly features. Treating cross-modal semantic features as the prior, a Semantic Guided Adapter (SGA) is designed to encode effective guidance signals for the adequate and controllable synthesis process. Extensive experiments indicate that AnomalyControl can achieve state-of-the-art results in anomaly synthesis compared with existing methods while exhibiting superior performance for downstream tasks.
@@
Figure 1: Learning cross-modal semantic features as the prior, our method can achieve both realism and generalization for controllable anomaly synthesis. Realism: The targeted text prompt provides a full description for the targeted anomaly image to be generated (e.g., color, texture, and shape), while the text-image reference prompt provides a template reference for the targeted anomaly region (e.g., magnified area). Generalization: Our method does not require the matched pair of text-image reference prompt and targeted text prompt, e.g., the text-image reference can exhibit a different material surface compared with the targeted text prompt (e.g., the right examples).
1 Introduction
@@
Anomaly inspection plays a crucial role in various fields, from industrial anomaly detection [17, 38] to medical imaging [14]. One of the most challenging issues is the scarcity of diverse abnormal data due to the resource-intensive procedure for acquiring large and varied datasets. Therefore, anomaly synthesis task [34, 7, 39, 22, 8, 13, 12, 35, 21] gradually emerged as an advancing technique to improve the data scale of the available abnormal sample.
@@
Many downstream tasks, such as anomaly detection and localization, could benefit from enriching the training dataset using synthetic anomaly samples. Recently, text-to-image synthesis techniques [26, 31] have received more research attention for their outperforming performance, which can generate visually realistic natural images according to the input text prompt. However, for the anomaly synthesis task, they still suffer from insufficient descriptors of the textual prompt, resulting in unsatisfactory realism for the generated anomalies. This limitation hampers the model’s ability to recognize rare or previously unseen defect types.
@@
Motivated by the success of publicly available large text-to-image diffusion models [29, 27, 20, 2, 32], many approaches proposed to incorporate extra control signals to assist the diffusion process, which could enrich the targeted text prompt via injecting more prior information. For example, ControlNet [37], Uni-ControlNet [40], and IP-Adapter [33] proposed to apply adapters on stable diffusion models [27] to provide extra control for generation. The main idea is to learn stylistic information from a reference image (called image prompt). However, image prompts without a clear focal point generally offer weak alignment signals, limiting the effectiveness of generating high-fidelity and customized images. This limitation is especially pronounced in anomaly synthesis tasks, where anomalies typically occupy only a small region of the image and can easily be overlooked during generation. As a result, the synthesized anomalies often lack sufficient details to match real samples, failing to meet the high precision required for effective anomaly inspection and resulting in a lack of realism in the synthesized samples.
@@
To this end, this work aims to enhance the controllability of anomaly synthesis, improving the realism of synthesized samples and supporting more flexible and generalizable prompts. The resulting framework is called AnomalyControl (as shown in Figure 2). Specifically, the controllability signals come from a non-matching pair, i.e., a text-image reference prompt and a targeted text prompt. The text-image reference prompt includes a textual anomaly descriptor and visual anomaly descriptor, which are used to offer semantic and contextual cues for guiding the synthesis process. Importantly, the anomaly described by the text-image reference prompt is aligned with that indicated in the targeted text prompt, while the background and other properties are left unrestricted, allowing for greater flexibility in anomaly placement. To improve realism, we introduce Cross-modal Semantic Modeling (CSM) to capture cross-modal semantic features via multimodal interactions, where the Anomaly-Semantic Enhanced Attention (ASEA) mechanism could extract fine-grained information related to the anomaly region. By taking cross-modal semantic features as the prior, the Semantic Guided Adapter (SGA) could insert effective semantic signals to support an adequate and controllable synthesis process. Table 1 makes a comparison between AnomalyControl and previous methods, where our method exhibits several advantages in terms of realism, generalization, plug-and-play functionality, and controllability.
@@
  • A novel framework called AnomalyControl is proposed to achieve controllable anomaly synthesis in a flexible manner, which can generate realistic and generalized anomaly samples guided by a non-matching pair.
  • We design a Cross-modal Semantic Modeling (CSM) module to extract cross-modal semantic features from the textual and visual anomaly descriptors, where an Anomaly-Semantic Enhanced Attention (ASEA) mechanism can encode fine-grained visual details into cross-modal semantic features via multimodal interactions, which could further improve the realism of generated anomalies.
  • We formulate a Semantic Guided Adapter (SGA) allowing cross-modal semantic features to pass through the generation procedure (i.e., diffusion process), providing adequate and controllable synthesis instructions to generate high-quality anomalies.
  • Extensive experiments demonstrate that our method can generate realistic and generalized anomalies based on different control signals (as shown in Figure 1) and significantly enhance anomaly inspection performance.
Table 1: Comparisons with previous methods. Realism refers to the ability to generate realistic anomaly images with complex and fine-grained details. Generalization indicates the capability to transfer anomaly features across different objects and surfaces, allowing the model to generate diverse anomalies beyond the knowledge learned from pre-training. Anomaly synthesis methods could prioritize high-fidelity anomaly data for downstream tasks, where the generalization ability is limited by the knowledge learned from pre-training. Benefiting from large-scale pre-training, image generation methods dominate the slightly stronger generalization ability for natural images, but they still suffer from low-quality anomaly synthesis due to the complicated characteristics of anomalies. Our AnomalyControl can perform well for realism and generalization, along with plug-and-play flexibility and controllability.
@@
DRAEM [34] ICCV’21 Anomaly Synthesis
DefGAN [35] WACV’21 Anomaly Synthesis
DFMGAN [8] AAAI’23 Anomaly Synthesis
AnoDiff [13] AAAI’24 Anomaly Synthesis
AnoXFusion [12] Arxiv’24 Anomaly Synthesis
ControlNet [37] ICCV’23 Image Generation
Uni-ControlNet [40] NeurIPS’23 Image Generation
IP-Adapter [33] Arxiv’23 Image Generation
@@
2 Related Work
2.1 Anomaly Synthesis
@@
Anomaly synthesis has become an essential technique to support anomaly detection systems, especially in scenarios where real defect data is scarce. Existing methods can be broadly divided into two main categories. First, crop-and-paste anomaly augmentation methods [34, 7] combine normal images with abnormal patterns sourced from small abnormal samples or external textures. While these methods are computationally efficient and straightforward to implement, they often lack realism and generalization, limiting their ability to represent the complex characteristics of anomalies. Second, generative model-based methods [39, 22, 8, 13, 12, 35, 21], such as GANs and diffusion models, generate anomalies from scratch or by fine-tuning pre-trained models. Although these methods typically perform well for in-distribution anomalies, they require sufficient sample generalization and frequently struggle to produce out-of-distribution anomalies, which limits their ability to generalize to unseen defect types.
2.2 Text-to-Image Diffusion Models
@@
Since fine-tuning large pre-trained models is often inefficient, using adapters has become a more efficient alternative. Adapters introduce a small number of trainable parameters into the original model while freezing the pre-trained weights, significantly reducing training costs. With the success of large text-to-image diffusion models like DALL-E 2 [26], Imagen [29], and Stable Diffusion [27], significant progress has been made in image generation. However, writing complex text prompts remains a challenge, and pure text often fails to fully express complex scenes or concepts. To address this, methods such as ControlNet [37] and Uni-ControlNet [40] have introduced lightweight adapters in pre-trained Stable Diffusion models, enabling more precise control by incorporating additional prompts, such as structural or image-based control. IP-Adapter [33] further optimizes this approach by introducing a decoupled cross-attention mechanism, improving the effectiveness of image prompting, and even achieving performance comparable to fine-tuning the entire model. However, image prompts without a clear focus often provide weak alignment signals, making it difficult to generate high-fidelity, customized images, especially in anomaly synthesis, where anomalies are small and easily overlooked.
Figure 2: Overall pipeline of the proposed AnomalyControl. Our framework consists of three main modules to enhance controllability in anomaly synthesis. (1) Cross-modal Semantic Modeling (CSM) integrates both visual anomaly descriptors $I_{a}$ and textual anomaly descriptors $T_{a}$ to capture precise cross-modal semantic features using a frozen vision-language model (VLM). (2)Anomaly-Semantic Enhanced Attention (ASEA) employ a trainable attention guidance variable $e_{g}$ to emphasize the designated anomaly regions, which can capture more accurate semantic information without retraining the frozen VLM. (3) Semantic Guided Adapter (SGA) utilizes a decoupled cross-attention to support controllable anomaly synthesis, allowing for a flexible and generalized generation.
3 Methods
3.1 Preliminaries
@@
Stable Diffusion. Our method is built upon Stable Diffusion [27], which efficiently performs the diffusion process in a low-dimensional latent space instead of pixel space by using an auto-encoder [15]. Specifically, given an input image $x_{i}\in\mathbb{R}^{H\times W\times 3}$, the encoder maps it to a latent representation $z_{0}=\xi(x_{i})$, where $z_{0}\in\mathbb{R}^{h\times w\times c}$ with $H/h=W/w$ as the downsampling factor and $c$ as the number of latent dimensions. The diffusion process utilizes a denoising UNet [28] to gradually refine a noisy latent $z_{t}$, conditioned on the current timestep $T$ and a textual prompt embedding $C$. Here, $C$ represents the embedding of the targeted text prompt, generated by a pre-trained CLIP [25] text encoder. The training objective is defined as:
@@
Overall Framework. The pipeline of AnomalyControl is illustrated in Figure 2. Our framework consists of two main modules: Cross-modal Semantic Modeling (CSM) and Semantic Guided Adapter (SGA). Given the text-image reference prompts as the input, i.e., a visual anomaly descriptor $I_{a}$ and textual anomaly descriptor $T_{a}$ as the reference, CSM module aims to extract fine-grained cross-modal semantic features by integrating both text and image information using a vision-language model (VLM). An Anomaly-Semantic Enhanced Attention (ASEA) is introduced to refine the attention process in VLM, ensuring that the model can focus on the specific visual patterns of the anomaly, thus enhancing the realism and contextual relevance of the generated anomaly features. The SGA then introduces these enriched cross-modal semantic features as priors into the diffusion process, allowing flexible control over anomaly characteristics. Together, these components ensure that generated anomalies are realistic and contextually aligned with descriptive cues. The detailed training algorithm of our AnomalyControl is outlined in Algorithm 1.
3.2 AnomalyControl
3.2.1 Cross-modal Semantic Modeling
Leveraging the text-image reference prompt, the CSM module targets to encode sufficient cross-modal reference information as control signals for the diffusion process, i.e., consistent anomaly semantics from the text-image prompt.
@@ 3.2.1 Cross-modal Semantic Modeling
Specifically, the text-image reference prompt provides a visual anomaly descriptor $I_{a}$ and a textual anomaly descriptor $T_{a}$. Note that such text-image reference prompt only focuses on the contextual and descriptive information of the anomaly region, excluding the anomaly-unrelated information, such as background and materials. Such a pattern would allow greater flexibility in anomaly placement. In detail, $I_{a}$ defines a specific anomaly pattern to provide a direct visual context reference, ensuring accurate visual control over the generated anomaly. $T_{a}$ is created by combining an anomaly-specific keyword $K$ with a template, providing semantic guidance to help the model generate samples that accurately reflect the designated anomaly type and attributes. The anomaly described in the $T_{a}$ aligns with the anomaly region in the $I_{a}$, while the background and other elements of $I_{a}$ remain unconstrained.
@@
To capture effective cross-modal information from the text-image reference, a frozen VLM with cross-attention layers is utilized for its powerful multimodal integration ability. Thus, VLM-based CSM could integrate the reference image and text to obtain a unified semantic feature for the anomaly, called a cross-modal semantic feature, which serves as a prior for the following diffusion process. Natural needs arise from the anomaly-specific regions without losing fine-grained detail to generate realistic and contextually accurate anomalies, which motivates the introduction of the ASEA mechanism in the following section.
3.2.2 Anomaly-Semantic Enhanced Attention
@@
To further refine the attention on anomaly regions, we apply an anomaly region mask $M_{A}$ to $\bar{A}_{\text{anomaly}}$. This mask constrains attention to the spatial area of interest in the image, which is enhanced by introducing the trainable attention guidance variable $e_{g}$. The optimization of $e_{g}$ is guided by an energy function [6], defined as:
@@
3.2.3 Semantic Guided Adapter
Motivated by the decoupled cross-attention [33], the SGA module introduces an adaptive mechanism to incorporate an additional targeted text prompt $T_{t}$ as input, indicating the specific anomaly to be generated. Unlike previous approaches using solely image features as the prior [33, 40], which often miss critical details in subtle anomaly regions with low signal-to-noise ratios, SGA relies on a richer, semantically focused representation, i.e., cross-modal semantic features produced by CSM, which captures fine-grained anomaly details essential for accurate synthesis.
To facilitate classifier-free guidance [11], similar to text conditioning, SGA employs random dropout during training. This process involves occasionally setting cross-modal semantic features to zero, allowing the model to jointly learn both conditional and unconditional prompts. The enhanced cross-attention mechanism for integrating text and cross-modal semantic features is defined as:
@@ 3.2.3 Semantic Guided Adapter
where $\gamma$ is a weight to balance these two terms. $\mathcal{S}$ is the function of Softmax. $Q$, $K$, and $V$ are the query, key, and value matrices for the attention operation applied to text cross-attention, while $K^{\prime}$ and $V^{\prime}$ correspond to cross-modal attention. Given the query features $Z$ and cross-modal semantic features $C^{\prime}$, the query matrix $Q$ is defined as $Q=ZW_{q}$, with $K^{\prime}=C^{\prime}W_{k}^{\prime}$ and $V^{\prime}=C^{\prime}W_{v}^{\prime}$. Notably, only $W_{k}^{\prime}$ and $W_{v}^{\prime}$ are trainable parameters, focusing the adaptation on cross-modal information. During training, only the parameters of the SGA and $e_{g}$ are optimized, while the pre-trained diffusion model and VLM parameters remain frozen. The entire AnomalyControl pipeline is trained on image-text pairs of anomaly images, with a training objective based on the original Stable Diffusion work:
@@
$\mathcal{L}=\mathbb{E}_{z_{t},t,C,C^{\prime},\epsilon\sim\mathcal{N}(0,1)}\left[\|\epsilon-\epsilon_{\theta}(z_{t},t,C,C^{\prime})\|^{2}_{2}\right].$ (6)

### v4

@@ Abstract
Anomaly synthesis is a crucial approach to augment abnormal data for advancing anomaly inspection. Benefiting from large-scale pre-training, existing text-to-image anomaly synthesis methods predominantly focus on textual information or coarse-aligned visual features to guide the entire generation process. However, these methods often lack sufficient descriptors to capture the complicated characteristics of realistic anomalies (e.g., the fine-grained visual pattern of anomalies), limiting the realism and generalization of the generation process. To this end, we propose a novel anomaly synthesis framework called AnomalyControl to learn cross-modal semantic features as guidance signals, which could encode the generalized anomaly cues from text-image reference prompts and improve the realism of synthesized abnormal samples. Specifically, AnomalyControl adopts a flexible and non-matching prompt pair (i.e., a text-image reference prompt and a targeted text prompt)  [^1] Non-matching: the text-image reference prompt and targeted text prompt are only required to describe the same anomaly type without needing to match in surface details or materials., where a Cross-modal Semantic Modeling (CSM) module is designed to extract cross-modal semantic features from the textual and visual descriptors. Then, an Anomaly-Semantic Enhanced Attention (ASEA) mechanism is formulated to allow CSM to focus on the specific visual patterns of the anomaly, thus enhancing the realism and contextual relevance of the generated anomaly features. Treating cross-modal semantic features as the prior, a Semantic Guided Adapter (SGA) is designed to encode effective guidance signals for the adequate and controllable synthesis process. Extensive experiments indicate that AnomalyControl can achieve state-of-the-art results in anomaly synthesis compared with existing methods while exhibiting superior performance for downstream tasks.
Codehttps://github.com/DaniellaHe/AnomalyControl
Introduction
@@
Figure 1: Our method learns cross-modal semantic features as the priors to achieve both realism and generalization for controllable anomaly synthesis. Realism: The targeted text prompt provides a full description for the targeted anomaly image (e.g., color, texture, and shape), while the text-image reference prompt provides a template for the targeted anomaly region (e.g., magnified area). Generalization: The text-image reference can exhibit a different material surface compared with the targeted text prompt (e.g., the right examples).
@@
Anomaly inspection plays a crucial role in various fields, from industrial anomaly detection (Liu et al. 2024; Zhang, Xu, and Zhou 2024) to medical imaging (Huang et al. 2024). One of the most challenging issues is the scarcity of diverse abnormal data due to the resource-intensive procedure for acquiring large and varied datasets. Therefore, anomaly synthesis task (Zavrtanik, Kristan, and Skočaj 2021; Chen et al. 2023; Zhao et al. 2020; Ojha et al. 2021a; Duan et al. 2023; Hu et al. 2024b, a; Zhang et al. 2021; Niu et al. 2020; Gui et al. 2024) gradually emerged as an advancing technique to improve the data scale of the available abnormal sample.
@@
Many downstream tasks, such as anomaly detection and localization, could benefit from enriching the training dataset using synthetic anomaly samples. Recently, text-to-image synthesis techniques (Ramesh et al. 2022; Sun, Cao, and Fink 2024) have received more research attention for their outperforming performance, which can generate visually realistic natural images according to the input text prompt. However, for the anomaly synthesis task, they still suffer from insufficient descriptors of the textual prompt, resulting in unsatisfactory realism for the generated anomalies. This limitation hampers the model’s ability to recognize rare or previously unseen defect types.
@@
Motivated by the success of publicly available large text-to-image diffusion models (Saharia et al. 2022; Rombach et al. 2022; Nichol et al. 2021; Balaji et al. 2022; Xue et al. 2024), many approaches have been proposed to incorporate extra control signals to assist the diffusion process, which could enrich the targeted text prompt via injecting more prior information. For example, ControlNet (Zhang, Rao, and Agrawala 2023), Uni-ControlNet (Zhao et al. 2023), and IP-Adapter (Ye et al. 2023) proposed to apply adapters on stable diffusion models (Rombach et al. 2022) to provide extra control for generation. The main idea is to learn stylistic information from a reference image (called an image prompt). However, image prompts without a clear focal point generally offer weak alignment signals, limiting the effectiveness of generating high-fidelity and customized images. This limitation is especially pronounced in anomaly synthesis tasks, where anomalies typically occupy only a small region of the image and can easily be overlooked during generation. As a result, the synthesized anomalies often lack sufficient details to match real samples, failing to meet the high precision required for effective anomaly inspection and resulting in a lack of realism in the synthesized samples.
@@
To this end, this work aims to improve the realism of synthesized samples via enhanced controllability of anomaly synthesis. The resulting framework is called AnomalyControl (as shown in Figure 2), supporting more flexible and generalizable prompts. Specifically, the controllability signals come from a non-matching pair, i.e., a text-image reference prompt, and a targeted text prompt. The text-image reference prompt includes both textual and visual anomaly descriptors to offer semantic and contextual cues for the synthesis process. To improve realism, we introduce Cross-modal Semantic Modeling (CSM) to capture cross-modal semantic features via multimodal interactions, where the Anomaly-Semantic Enhanced Attention (ASEA) mechanism could extract fine-grained information related to the anomaly region. By taking cross-modal semantic features as the prior, the Semantic Guided Adapter (SGA) could insert effective semantic signals to support an adequate and controllable synthesis process. Table 1 makes a comparison between AnomalyControl and previous methods, where our method exhibits several advantages in four aspects: 1) Realism: Our model generates highly realistic anomalies by using cross-modal semantic features, closely matching real-world defects. 2) Generalization: Synthesizes diverse anomalies beyond the training distribution, improving adaptability to unseen cases. 3) Plug-and-Play: The Semantic Guided Adapter (SGA) allows flexible anomaly synthesis without retraining the diffusion model, making it efficient and scalable. 4) Controllability: AnomalyControl provides precise control over anomaly types and placements using text-image reference and targeted text prompts.
@@
(1) A novel framework called AnomalyControl is proposed to achieve controllable anomaly synthesis in a flexible manner, which can generate realistic and generalized anomaly samples guided by a non-matching pair.
(2) We design a Cross-modal Semantic Modeling (CSM) module to extract cross-modal semantic features from the textual and visual anomaly descriptors, where an Anomaly-Semantic Enhanced Attention (ASEA) mechanism can encode fine-grained visual details into cross-modal semantic features via multimodal interactions, which could further improve the realism of generated anomalies.
(3) We formulate a Semantic Guided Adapter (SGA) allowing cross-modal semantic features to pass through the generation procedure (i.e., diffusion process), providing adequate and controllable synthesis instructions to generate high-quality anomalies.
(4) Extensive experiments demonstrate that our method can generate realistic and generalized anomalies based on different control signals (as shown in Figure 1) and significantly enhance anomaly inspection performance.
@@
DRAEM (Zavrtanik, Kristan, and Skočaj 2021) ICCV’21 Anomaly Synthesis
DefGAN (Zhang et al. 2021) WACV’21 Anomaly Synthesis
DFMGAN (Duan et al. 2023) AAAI’23 Anomaly Synthesis
AnoDiff (Hu et al. 2024b) AAAI’24 Anomaly Synthesis
ControlNet (Zhang, Rao, and Agrawala 2023) ICCV’23 Image Generation
Uni-ControlNet (Zhao et al. 2023) NeurIPS’23 Image Generation
IP-Adapter (Ye et al. 2023) Arxiv’23 Image Generation
@@
Table 1: Comparisons with previous methods. Realism refers to the ability to generate realistic anomaly images with complex and fine-grained details. Generalization indicates the capability to transfer anomaly features across different objects and surfaces, allowing the model to generate diverse anomalies beyond the pre-training knowledge. Our AnomalyControl can perform well for realism and generalization, along with plug-and-play flexibility and controllability.
Related Work
@@
Anomaly Synthesis. Anomaly synthesis has become an essential technique to support anomaly detection systems, especially in scenarios where real defect data is scarce. Existing methods can be broadly divided into two main categories. First, crop-and-paste anomaly augmentation methods (Zavrtanik, Kristan, and Skočaj 2021; Chen et al. 2023) combine normal images with abnormal patterns sourced from small abnormal samples or external textures. While these methods are computationally efficient and straightforward to implement, they often lack realism and generalization, limiting their ability to represent the complex characteristics of anomalies. Second, generative model-based methods (Zhao et al. 2020; Ojha et al. 2021a; Duan et al. 2023; Hu et al. 2024b, a; Gui et al. 2024; Zhang et al. 2021; Niu et al. 2020), such as GANs and diffusion models, generate anomalies from scratch or by fine-tuning pre-trained models. Although these methods typically perform well for in-distribution anomalies, they require sufficient sample generalization and frequently struggle to produce out-of-distribution anomalies, which limits their ability to generalize to unseen defect types.
@@
Text-to-Image Diffusion Models. Since fine-tuning large pre-trained models is often inefficient, using adapters has become a more efficient alternative. Adapters introduce a small number of trainable parameters into the original model while freezing the pre-trained weights, significantly reducing training costs. With the success of large text-to-image diffusion models like DALL-E 2 (Ramesh et al. 2022), Imagen (Saharia et al. 2022), and Stable Diffusion (Rombach et al. 2022), significant progress has been made in image generation. However, writing complex text prompts remains a challenge, and pure text often fails to fully express complex scenes or concepts. To address this, methods such as ControlNet (Zhang, Rao, and Agrawala 2023) and Uni-ControlNet (Zhao et al. 2023) have introduced lightweight adapters in pre-trained Stable Diffusion models, enabling more precise control by incorporating additional prompts, such as structural or image-based control. IP-Adapter (Ye et al. 2023) further optimizes this approach by introducing a decoupled cross-attention mechanism, improving the effectiveness of image prompting, and even achieving performance comparable to fine-tuning the entire model. However, image prompts without a clear focus often provide weak alignment signals, making it difficult to generate high-fidelity, customized images, especially in anomaly synthesis, where anomalies are small and easily overlooked.
Methods
@@
Preliminaries. Our method is built upon Stable Diffusion (Rombach et al. 2022), which efficiently performs the diffusion process in a low-dimensional latent space instead of pixel space using an auto-encoder (Kingma 2013). Specifically, given an input $x_{i}\in\mathbb{R}^{H\times W\times 3}$, the encoder maps it to a latent representation $z_{0}=\xi(x_{i})$, where $z_{0}\in\mathbb{R}^{h\times w\times c}$ with $H/h=W/w$ as the downsampling factor and $c$ as the number of latent dimensions. The diffusion process utilizes a denoising UNet (Ronneberger, Fischer, and Brox 2015) to gradually refine a noisy latent $z_{t}$, conditioned on the current timestep $T$ and a textual prompt embedding $C$. $C$ represents the embedding of the targeted text prompt, generated by a pre-trained CLIP (Radford et al. 2021) text encoder. The training objective is defined as:
@@
Figure 2: Pipeline of the proposed AnomalyControl. Our framework consists of three modules to enhance controllability in anomaly synthesis. (1) CSM integrates both visual and textual anomaly descriptors ($I_{a}$, $T_{a}$) to capture precise cross-modal semantic features using a frozen VLM. (2) ASEA employs a trainable attention guidance variable $e_{g}$ to emphasize the designated anomaly regions, which can capture more accurate semantic information without retraining the frozen VLM. (3) SGA utilizes a decoupled cross-attention to support controllable anomaly synthesis, allowing for a flexible and generalized generation.
@@
Overall Framework. The pipeline of AnomalyControl is illustrated in Figure 2. Our framework consists of two main modules: Cross-modal Semantic Modeling (CSM) and Semantic Guided Adapter (SGA). Given the text-image reference prompts as the input, i.e., a visual anomaly descriptor $I_{a}$ and textual anomaly descriptor $T_{a}$ as the reference, the CSM module aims to extract fine-grained cross-modal semantic features by integrating both text and image information using a vision-language model (VLM). An Anomaly-Semantic Enhanced Attention (ASEA) is introduced to refine the attention process in VLM, ensuring that the model can focus on the specific visual patterns of the anomaly, thus enhancing the realism and contextual relevance of the generated anomaly features. The SGA then introduces these enriched cross-modal semantic features as priors into the diffusion process, allowing flexible control over anomaly characteristics. Together, these components ensure that generated anomalies are realistic and contextually aligned with descriptive cues.
AnomalyControl
Cross-modal Semantic Modeling
@@
Leveraging the text-image reference prompt, the CSM module targets to encode sufficient cross-modal reference information as control signals for the diffusion process, i.e., consistent anomaly semantics from the text-image prompt. Specifically, the text-image reference prompt provides a visual anomaly descriptor $I_{a}$ and a textual anomaly descriptor $T_{a}$. Note that such a text-image reference prompt only focuses on the contextual and descriptive information of the anomaly region, excluding the anomaly-unrelated information, such as background and materials. Such a pattern would allow greater flexibility in anomaly placement. In detail, $I_{a}$ defines a specific anomaly pattern to provide a direct visual context reference, ensuring accurate visual control over the generated anomaly. $T_{a}$ is created by combining an anomaly-specific keyword $K$ with a template, providing semantic guidance to help the model generate samples that accurately reflect the designated anomaly type and attributes. The anomaly described in the $T_{a}$ aligns with the anomaly region in the $I_{a}$, while the background and other elements of $I_{a}$ remain unconstrained.
@@
To capture effective cross-modal information from the text-image reference, a frozen VLM with cross-attention layers is utilized for its powerful multimodal integration ability. VLM-based CSM could integrate the reference image and text to obtain a cross-modal unified semantic feature, which serves as a prior for the following diffusion process. Natural needs arise from the anomaly-specific regions without losing fine-grained detail to generate realistic and contextually accurate anomalies, which motivates the introduction of the ASEA mechanism in the following section.
Anomaly-Semantic Enhanced Attention
@@
To further refine the attention on anomaly regions, we apply an anomaly region mask $M_{A}$ to $\bar{A}_{\text{anomaly}}$. This mask constrains attention to the spatial area of interest in the image, which is enhanced by introducing the trainable attention guidance variable $e_{g}$. The optimization of $e_{g}$ is guided by an energy function (Chen, Laina, and Vedaldi 2024), defined as:
@@
DiffAug IS 1.59 1.72 1.34 1.19 1.96 1.67 2.07 1.58 1.53 1.10 1.93 1.33 1.34 2.05 1.30 1.58
CDC IS 1.52 1.97 1.37 1.25 1.97 1.97 1.80 1.55 1.56 1.13 2.10 1.63 1.61 2.05 1.30 1.65
SDGAN IS 1.57 1.89 1.49 1.18 1.95 1.85 2.04 1.45 1.61 1.17 2.53 1.78 1.76 2.12 1.25 1.71
DefGAN IS 1.39 1.70 1.59 1.24 2.01 1.87 2.12 1.47 1.61 1.19 2.35 1.85 1.47 2.19 1.25 1.69
DFMGAN IS 1.62 1.96 1.59 1.23 1.97 1.93 2.06 1.49 1.63 1.12 2.39 1.82 1.64 2.12 1.29 1.72
AnoDiff IS 1.58 2.13 1.59 1.16 2.04 2.13 1.94 1.96 1.61 1.28 2.54 1.68 1.57 2.33 1.39 1.80
Ours IS 1.63 2.14 1.69 1.18 2.26 2.12 2.08 1.98 1.64 1.25 2.54 1.80 1.66 2.20 1.40 1.84
Table 2: Quantitative results of generation evaluated by IS and IC-LPIPS (abbreviated as IL) on the MVTec AD dataset. Bold and underlined values indicate the best and second-best results, respectively.
Semantic Guided Adapter
Using solely image features as the prior (Ye et al. 2023; Zhao et al. 2023), previous approaches often miss critical details in subtle anomaly regions with low signal-to-noise ratios. Motivated by this, SGA relies on a richer, semantically focused representation, i.e., cross-modal semantic features produced by CSM, which captures fine-grained anomaly details essential for accurate synthesis. Concretely, based on the decoupled cross-attention (Ye et al. 2023), SGA introduces an adaptive mechanism to incorporate an additional targeted text prompt $T_{t}$ as input.
To facilitate classifier-free guidance (Ho and Salimans 2022), SGA employs random dropout during training. This process involves occasionally setting cross-modal semantic features to zero, allowing the model to jointly learn both conditional and unconditional prompts. The enhanced cross-attention mechanism for integrating text and cross-modal semantic features is defined as:
@@
where $\gamma$ is a weight to balance these two terms. $\mathcal{S}$ is the function of Softmax. $Q$, $K$, and $V$ are the query, key, and value matrices for the attention operation applied to text cross-attention, while $K^{\prime}$ and $V^{\prime}$ correspond to cross-modal attention. Given the query features $Z$ and cross-modal semantic features $F_{c}$, the query matrix $Q$ is defined as $Q=ZW_{q}$, with $K^{\prime}=F_{c}W_{k}^{\prime}$ and $V^{\prime}=F_{c}W_{v}^{\prime}$. Notably, only $W_{k}^{\prime}$ and $W_{v}^{\prime}$ are trainable parameters, focusing the adaptation on cross-modal information. During training, only the parameters of the SGA and $e_{g}$ are optimized, while the pre-trained diffusion model and VLM parameters remain frozen. The entire AnomalyControl pipeline is trained on image-text pairs of anomaly images, with a training objective based on the original Stable Diffusion work:
@@
$\mathcal{L}=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,1)}\left[\|\epsilon-\epsilon_{\theta}(\text{concat}(z_{t},M_{A},z_{m},),t,F_{t},F_{c})\|^{2}_{2}\right],$ (6)